. A bicycle manufacturer currently produces 298,000 units a year and expects output levels to remain steady in the future. It buys chains from an outside supplier at a price of $1.90 a chain.

The plant manager believes that it would be cheaper to make these chains rather than buy them. Direct in-house production costs are estimated to be only $1.50 per chain. The necessary machinery would cost $292,000 and would be obsolete after 10 years. This investment could be depreciated to zero for tax purposes using a 10-year straight-line depreciation schedule. The

plant manager estimates that the operation would require $31,000 of inventory and other working capital upfront (year 0), but argues that this sum can be ignored since it is recoverable at the end of the 10 years. Expected proceeds from scrapping the machinery after 10 years are

$21,900. If the company pays tax at a rate of 35% and the opportunity cost of capital is 15%, what is the net present value of the decision to produce the chains in-house instead of purchasing them from the supplier?
Answer #1
1] Cost of the machinery $       2,92,000
Increase in NWC $ 31,000
Initial outlay $       3,23,000
2] Savings in cost of purchase = 298000*1.9 = $       5,66,200
Less: Direct in house production costs = 298000*1.5 = $       4,47,000
Less: Depreciation = 292000/10 = $           29,200
Equals: Incremental NOI $           90,000
Less: Tax at 35% $           31,500
Equals: Incremental NOPAT $           58,500
Add: Depreciation $           29,200
Equals: Incremental annual OCF $           87,700
3] Recovery of NWC $           31,000
After tax salvage value = 21900*(1-35%) = $           14,235
Terminal non operating cash flows $           45,235
4] PV of incremental annual OCF = 87700*(1.15^10-1)/(0.15*1.15^10) = $       4,40,146
PV of terminal non operating cash flows = 45235/1.15^10 = $ 11,181
PV of cash inflows $       4,51,327
Less: Initial outlay $       3,23,000
NPV $       1,28,327
5] As the NPV of the decision to produce is positive, the firm
should start producing the chains in house.